OpenStudy (anonymous):

Express the given quantity as a single logarithm. ln(1+x^2)+2/1lnx−lnsinx

OpenStudy (tkhunny):

That doesn't quite make sense. Is the '1' in front of the ln(x) doing something?

OpenStudy (anonymous):

I am not sure this is how it shows..

OpenStudy (tkhunny):

I'm troubled with that. Maybe just ignore it. ln(a) + ln(b) = ln(a*b) ln(a) - ln(b) = ln(a/b) What else do you need?

OpenStudy (anonymous):

what do you mean

OpenStudy (tkhunny):

I mean that you should apply those principles to your original problem statement.

Directrix (directrix):

Is the expression [ln(1+x^2)+2] /[1ln (x)−ln(sinx)] what you intended when you posted ln(1+x^2)+2/1lnx−lnsinx ?

OpenStudy (anonymous):

yes i made the confusion.

OpenStudy (anonymous):